OpenStudy (anonymous):

Solve. Justify each step. Check the solution. 3x + 2 - x + 7 = 16

OpenStudy (anonymous):

3x+2-x+7=16 -2 -2 3x-x+7=14 +x +x -7 -7 4x=7 4 4 x=1.8

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

ur welcome